You have a 2.20-L balloon full of air at 23°C. To what Celsius temperature would you have to heat the balloon to double its volume if the pressure remained unchanged?
Answer #1

Temperature, T = 23o C = (23 +273.15) K = 296.15 K

Now PV = nRT which gives T/V = P/(nR) = constant for fixed P and n.

So we get T1/V1 = T2/V2

Putting the T1=296.15 K, V2 = 2×V1 in above equation we have :

296.15K/V1 = T2/(2V1) which gives T2 = 2×296.15 K = 592.3 K = (592.3- 273.15) oC = 319.15 o C.

So we would have to heat the balloon to 319.15 oC.

Comment in case of any doubt.